### v1

# Evaluating End-User Device Energy Models in Sustainability Reporting of Browser-Based Web Services

Abstract.
Sustainability reporting in web-based services increasingly relies on simplified energy and carbon models such as the Danish Agency of Digital Government’s Digst framework and the United Kingdom-based DIMPACT model. Although these models are widely adopted, their accuracy and precision remain underexplored. This paper presents an empirical study evaluating how well such models reflect actual energy consumption during realistic user interactions with common website categories. Energy use was measured across shopping, booking, navigation, and news services using predefined user flows executed on four laptop platforms. The results show that the commonly applied constant-power approximation ($P\times t$) can diverge substantially from measured energy, depending on website category, device type, and task characteristics. The findings demonstrate that model deviations are systematic rather than random and highlight the need for category-aware and device-reflective power parameters in reproducible sustainability reporting frameworks.
Energy Consumption, End-user device, Energy Models, Reporting
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The 41st ACM/SIGAPP Symposium on Applied Computing; March 23–27, 2026; Thessaloniki, Greeceisbn: 979-X-XXXX-XXXX-X/26/03ccs: General and reference Evaluationccs: Software and its engineering Empirical software validation
1. Introduction
Digital sustainability reporting has become central to the evaluation of web-based services. Governments and organizations increasingly depend on simplified models to estimate energy use and greenhouse gas emissions. Two such models are the Danish Agency of Digital Government’s Digst framework (digst, 7) and the UK-based DIMPACT framework (dimpact, 8, 26) based on (Preist2019, 24). However, despite their growing role, the empirical validation of these models is limited. Without systematic assessment, it remains unclear how accurately they capture the energy demands of real-world web interactions.
This paper addresses this gap by presenting an experiment that directly compares measured energy consumption against model-based estimates. Following the recommendation by Guldner et al. (Guldner2024, 15) to evaluate software energy efficiency under representative user scenarios, and inspired by the online-media categories proposed by Schien et al. (Schien2013, 27)—who modeled energy use across text-based, interactive, and streaming web content—we focus on non-streaming web services (shopping, booking, navigation, and news). In doing so, we complement prior work that has predominantly investigated streaming services. While categories such as streaming and generative AI are highly relevant to digital sustainability, their energy use depends on complex scaling parameters (e.g., video quality (Suski2020, 28) or token length) that require separate, parameter-specific studies. We therefore delimit our study to web services with more stable and comparable interaction patterns.
Existing sustainability reporting frameworks for web-based services, such as Digst and DIMPACT, estimate end-user device energy use through constant-power assumptions. While these models enable comparability across services, their empirical accuracy and the factors influencing their validity remain largely unexplored. To address this gap, we investigate two research questions:
  • RQ1:
    How accurately do current reporting models estimate the energy consumption of end-user devices during realistic web-service interactions?
  • RQ2:
    Which factors—such as device type, website category, or interaction duration—explain systematic deviations between modeled and measured energy use?
Our contributions are threefold: (1) a reproducible experimental methodology for measuring the energy use of web-service interactions, (2) empirical evidence on the accuracy and precision of Digst- and DIMPACT-style energy estimates, and (3) insights for improving fairness and transparency in digital sustainability reporting.
2. Background and Related Work
This section situates the study within existing research on the estimation and reporting of ICT-related energy use. We first review sector-level and national models that underpin current digital sustainability frameworks, focusing on their assumptions about end-user devices. We then outline software- and measurement-level approaches that aim to quantify energy use more precisely at the program or interaction level. Together, these strands of work frame the methodological gap that this paper addresses: linking reproducible, software-level energy measurements to the simplified models used in large-scale sustainability reporting.
2.1. Sector-Level Models
Global and sector-level ICT footprint models establish the broader context for digital sustainability research. Early studies, e.g., (Andrae2015, 1, 21, 2), raised awareness of the growing energy challenges, while Freitag et al. (Freitag2021, 11) analyzed the methodological assumptions that explain divergent estimates of Internet energy intensity. Later studies, such as Malmodin et al. (malmodin2024, 20), provide comprehensive assessments of the ICT sector’s total energy use and greenhouse gas emissions, emphasizing harmonized system boundaries across data centers, networks, and end-user devices.
These models typically employ top-down allocation and averaged intensity factors to distribute total sector energy across subdomains. As a result, they inherently contain uncertainties, and Malmodin et al. specifically acknowledge that “the number of user devices in operation and their average electricity consumption remain a more uncertain part in line with previous studies” (malmodin2024, 20). This uncertainty is echoed by Furberg et al. (Furberg2026, 12), who note that “to estimate the electricity use of devices, data on actual device usage patterns (e.g., operational service time per year) are required. In practice, studies on the ICT sector’s direct climate impact usually do not collect device usage data but rely on secondary data from various sources.”e.g., averages over multiple sources (malmodin2024, 20).
Although these studies establish the overall scale of ICT’s environmental impact, their level of aggregation prevents validation of individual services or user interactions. Consequently, sector models cannot determine whether specific web services, user actions, or design choices meaningfully reduce energy demand. These limitations motivate the development of service-level models for digital sustainability reporting, which attempt to translate aggregate insights into interpretable metrics for individual services.
2.2. Service-Level Models
Service-level models aim to make the environmental assessment of digital services practical and comparable. Frameworks such as the Danish Agency of Digital Government [^1] https://en.digst.dk/’s Digst model (digst, 7) and the United Kingdom-based DIMPACT tool (dimpact, 8, 26) operationalize simplified estimation methods that express energy use as the product of assumed device power and activity duration ($P\times t$). These models allocate a share of data-center and network energy to each service and are increasingly used for organizational sustainability reporting. The approach aligns with academic efforts such as Preist et al. (Preist2019, 24), who analyzed how interface design in YouTube affects energy demand, and with the Green Software Foundation’s Software Carbon Intensity (SCI) specification, which formalizes functional units and allocation principles for shared infrastructure.
Although these frameworks facilitate transparency and comparability, they rely on several simplifying assumptions: constant device power, uniform user behavior, and linear relationships between time, data volume, and energy use.
The precision and accuracy of these models have not yet been evaluated against measured data from real user interactions. Our study directly addresses this gap by empirically testing the validity of the $P\times t$ assumption across realistic user flows and hardware configurations. By comparing measured energy data with estimates derived from Digst- and DIMPACT-style models, we provide an evidence-based assessment of their accuracy and identify conditions under which these models deviate from actual energy consumption.
While these models target sustainability reporting, a parallel line of work addresses developer-facing concerns by improving software energy efficiency and promoting reproducible energy-measurement practices.
2.3. Software-Level Efficiency and Measurement Frameworks
Research on software-level energy efficiency has produced a range of tools and methodologies for analyzing, predicting, and improving the energy behavior of programs. Early frameworks such as EACOF (Field2014, 10) enabled applications to access power data at runtime, fostering awareness of energy use during development. Static and data-dependent analyses by Grech et al. (Grech2015, 14) and Pallister et al. (Pallister2017, 22) estimate energy consumption directly from program structure or input characteristics, while comparative studies such as Georgiou et al. (Georgiou2017, 13) and Pereira et al. (Pereira2021, 23) quantify how programming language choice influences energy efficiency. Complementary approaches examine reusable solutions for efficient design—such as energy-related software patterns (Rani2024, 25, 9)—and energy-aware debugging and testing frameworks for mobile and desktop applications.
More recently, Guldner et al. (Guldner2024, 15) proposed the Green Software Measurement Model, a standardized process for measuring software resource efficiency under representative user scenarios. Their rationale is that meaningful and reproducible results require measurements grounded in realistic user interactions rather than synthetic benchmarks. This principle directly aligns with our methodological choice to evaluate energy consumption through end-to-end user flows across web services.
While these works provide methodological foundations for measuring and improving software efficiency, their focus remains on optimizing or characterizing software behavior in isolation. Our study builds upon their measurement philosophy but applies it to a different objective: validating the accuracy of service-level reporting models on real-world services.
2.4. Web-Specific Energy Studies
Energy consumption in web-based services has been investigated from multiple perspectives, ranging from modeling of user-device interactions to experimental measurement of browser and video workloads. Schien et al. (Schien2013, 27) provided one of the earliest models of online media consumption, distinguishing between text-based, interactive, and streaming content across devices such as desktops, laptops, and mobile phones. Their results revealed that end-user devices, particularly laptops, dominate overall energy consumption, and that usage patterns substantially influence total impact. This modeling work established a foundation for categorizing web services by data intensity and interaction type—a principle we adopt when structuring our own experimental categories.
Subsequent studies examined specific behavioral and technical factors that shape web energy use. Suski et al. (Suski2020, 28) demonstrated that user decisions, such as video resolution and device choice, strongly affect streaming emissions. Rani et al. (Rani2024, 25) identified recurring energy-related design patterns in web applications, emphasizing that front-end implementation choices can meaningfully alter client-side power demand. Macedo et al. (Macedo2022, 3) compared WebAssembly and JavaScript, showing that runtime technology influences both execution time and energy efficiency. Together, these studies highlight the diversity of determinants—behavioral, architectural, and technological—that drive energy use in web contexts.
Recent work has advanced the methodological basis for measuring energy in multimedia and browser environments. Katsenou et al. (Katsenou2024, 18) compared hardware- and software-based power measurements in video compression, while Herglotz et al. (Herglotz2025, 16) formalized procedures for synchronized power and time measurements in video communication systems. Their controlled setups provide technical precision that complements our empirical approach. In our study, news-site user flows occasionally included embedded video playback, allowing us to relate these real-world observations to the controlled results reported by Herglotz et al. Similarly, our measured laptop data provide empirical confirmation of Schien et al.’s modeled dominance of device energy use.
Collectively, this literature demonstrates that web-service energy use is heterogeneous and context dependent. However, few of these studies examine whether the simplified assumptions used in service-level reporting models can capture such variability. Our work addresses this gap by providing measured, category-based data to test the validity of those assumptions under realistic web interactions.
2.5. Positioning of This Work
Existing literature thus ranges from sector-wide modeling to code-level analysis, with few studies bridging measured data and reporting models. Our study occupies this intermediate layer: it empirically tests the accuracy and precision of service-level energy models (Digst and DIMPACT) under realistic user flows. Following the recommendation of Guldner et al. (Guldner2024, 15) and building on the content categories proposed by Schien et al. (Schien2013, 27), we structure our experiment around four non-streaming web-service categories—shopping, booking, navigation, and news—and later relate our results to prior findings on device-level energy variation and video playback effects. The following section details our experimental design, which applies these principles in a controlled setting to test model validity across four representative web-service categories.
3. Methodology
This section describes the experimental design, data collection, and model evaluation procedures used to address the research questions. We outline the selection of representative web services and user flows, the hardware and participant setup, and the measurement process for recording power and duration. Finally, we detail how the measured data were compared with the simplified energy models used in existing reporting frameworks to assess their accuracy and structural bias.
3.1. Websites and User Flows
We selected eight popular websites grouped into four categories:
  1. (1)
    Shopping: ebay.fr, amazon.fr;
  2. (2)
    Booking: hotels.com, booking.com;
  3. (3)
    Directions: google.com/maps, bing.com/maps; and
  4. (4)
    News: dr.dk, tv2.dk
For each category, we defined realistic user flows that mimic everyday usage—searching for a product, reserving a hotel room, generating a route, or reading a news article—so that equivalent functionality was tested across services. Streaming and other parameter-dependent services were deliberately excluded to ensure comparability of interaction length and structure.
3.2. Devices
Experiments were performed on four laptops representing different hardware generations and OEM configurations:
  1. Laptop 1:
    Dell Latitude 7490 (Intel 8th-gen, UHD 620),
  2. Laptop 2:
    HP EliteBook 830 G6 (Intel 8th-gen, alternate firmware),
  3. Laptop 3:
    Dell Latitude E5250 (Intel 5th-gen, compact chassis), and
  4. Laptop 4:
    Dell Latitude E5550 (Intel 5th-gen, larger chassis).
This selection captures both processor-generation and design-level power variation.
3.3. Participants and Experimental Design
Ten participants completed the experiment; one participant (P10) contributed only two runs and was excluded from the mixed-effects analyses to ensure stable estimation of random effects. The remaining nine participants provided between eight and twenty-four valid runs each. The design crossed eight user flows with four laptops, and every flow–computer combination was executed by at least four and up to seven distinct participants, ensuring balanced coverage. To counter potential learning or fatigue effects while maintaining full coverage, the order of user flows and the allocation of laptops were randomized under constraints that ensured each flow–device combination was represented across participants. Each flow was repeated three times per device, yielding a total of $164$ valid observations.
3.4. Experimental Setup and Measurements
All laptops ran identical operating-system images (Ubuntu Desktop 24.04.3 LTS) with default power settings and brightness. The same Chromium browser version was used in incognito mode, and caches were cleared between runs. Each measurement recorded device power (W) via a Siglent SPD3303X-E programmable power supply and task duration (s), allowing energy to be computed as integrated power over time. The external power-supply measurements was user-synchronized via task start and stop events to ensure precise alignment between user interaction and power logging. The setup can be seen in Figure1. Key sources of potential variance included machine architecture, user execution behavior, and learning effects, which were addressed through randomization and replication. Potential influences of these factors and the sample size are discussed in Section 7.
Figure 1. The setup for the experiment with 4 Laptops, two Siglents and two data-collecting computers with user-controlled synchronization software.
3.5. Model Evaluation
To evaluate how well sector-level reporting models represent actual end-user device energy, we implemented both the Digst and DIMPACT formulations. In both frameworks, energy use is estimated as constant device power multiplied by interaction duration ($E_{\mathrm{model}}=P_{\mathrm{avg}}\times t$).
Digst uses (digst, 7, Table 1 ) a composite averaging over laptops (22 W), tower desktops (87 W), all-in-one PCs (81 W), and monitors (31 W) to derive an overall mean user-PC power of approximately 68.5 W, i.e., the average power ratings reported by Urban et al. (urban2017, 29). However, since we focus on laptops, we have extracted the laptop power of 22 W. In comparison, DIMPACT have in the case of browser-based services chosen a lower constant of 15 W for web-browsing scenarios (dimpact, 8, 26), i.e., Urban’s power ratings without workload adjustment (urban2017, 29). We will use these constants (22W and 15W) to evaluating energy models in Section 5.1.
4. Reproducibility
All anonymized datasets, user-flow descriptions, and analysis scripts are available in the open repository EUD-energy-model [^2] https://github.com/SustainableSoftware/EUD-energy-model. The repository includes all data, scripts, version numbers, and documentation needed to reproduce the study and the analyses in this study.
5. Results
Our analysis examined accuracy and precision across categories, machines, and flows. Initial findings suggest that while power multiplied by time estimates are broadly correlated with measured energy, deviations occur systematically. In particular, shopping showed higher variance, while navigation tasks on maps platforms tasks produced more consistent estimates across devices, e.g., visualized in Figures 7 and  8. Machine differences (5th vs. 8th gen CPUs, OEM design) also contributed significantly to deviations, e.g., visualized in Figure 6.
5.1. Model Accuracy and Error Growth (RQ1)
To assess how accurately the national reporting models reflect measured device energy, we compared the estimated energy consumption from the DIMPACT and Digst frameworks against empirical measurements from controlled web-service user flows. Both models implement a simplified $P_{\mathrm{avg}}\times t$ formulation, where energy use is approximated as constant device power multiplied by task duration. In the Digst framework, this constant is set to 22 W based on national reference data, whereas DIMPACT applies a lower value of 15 W for web-browsing scenarios. These fixed parameters provide a straightforward baseline for reporting but inherently assume uniform device behavior across users and workloads.
Figure 2. Measured versus modeled energy consumption for the DIMPACT framework. The red dashed line indicates perfect prediction ($y=x$).
Figure 3. Measured versus modeled energy consumption for the Digst framework. The stronger deviation from the reference line reflects systematic underestimation.
Figure 4. Relationship between task duration and absolute percentage error in DIMPACT energy estimates ($\beta=5.0\,\text{J s}^{-1},R^{2}=0.86$).
Figure 5. Relationship between task duration and absolute percentage error in Digst energy estimates ($\beta=11.8\,\text{J s}^{-1},R^{2}=0.96$).
5.1.1. Model Accuracy
Figure 2 compares measured energy with DIMPACT’s model estimates. The values are positively correlated, but the model systematically underestimates actual energy use for higher-energy tasks. Across all flows, the mean percentage error was $26.2\,\%$, the mean absolute percentage error (MAPE) was $28.7\,\%$, and the root-mean-square error (RMSE) was $25.5\,\text{J}$. The corresponding Digst comparison (Figure 3) revealed a stronger underestimation bias, with a mean error of $49.7\,\%$, MAPE of $49.7\,\%$, and RMSE of $38.1\,\text{J}$. While both models show moderate overall agreement with measured values, the deviation from the 1:11{:}1 reference line confirms that the constant-power assumption does not hold for realistic browser interactions.
5.1.2. Systematic Growth of Error with Duration.
To test whether model error accumulates with task length, we examined the relationship between task duration and absolute percentage error. For DIMPACT (Figure 4), a linear regression explained $86\,\%$ of the variance in absolute error ($R^{2}=0.86,F(1,165)=989.5,p<0.001$), with a slope of $5.0\pm 0.16\,\text{J s}^{-1}$. For Digst (Figure 5), the relationship was even stronger, with $96\,\%$ of the variance explained ($R^{2}=0.96,F(1,165)=3929,p<0.001$) and a slope of $11.8\pm 0.19\,\text{J s}^{-1}$. These results show that model deviations increase almost linearly with interaction length: longer user flows yield proportionally larger relative errors. The pattern demonstrates that small per-second inaccuracies in the $P_{\mathrm{avg}}\times t$ approximation accumulate over time, leading to systematic underestimation for extended browsing sessions. Together, these findings confirm that although Digst and DIMPACT capture the general order of magnitude of web-service energy use, neither provides accurate or duration-independent estimates under realistic usage conditions.
5.2. Factors Impacting Model Deviations (RQ2)
To determine which factors explain the deviations observed in RQ1, we incrementally extended the proportional $P\times t$ model with parameters representing hardware and web-service variability. All models were fitted without intercepts to preserve the physical relationship ($E_{\mathrm{model}}=P_{\mathrm{avg}}\times t$). Table 2 summarizes the progression in explanatory power.
5.2.1. Hardware and Category Effects.
The hardware-extended model ($E\sim t\!:\!C(\text{computerid})-1$) showed systematic differences in effective power between laptops ($R^{2}=0.960$), ranging from approximately 9 W to 12 W. These differences align with the measured distributions of average power across devices shown in Figure 6. Such hardware variation reflects both CPU generation and OEM design, and accounts for a substantial share of the total variance in energy use.
At the category level, user-flow differences explain an even larger portion of the remaining variance. Figure 7 illustrates that total energy consumption varies systematically between user flows, with news and navigation tasks exhibiting the highest medians. Corresponding differences in average device power (Figure 8) confirm that these effects stem from task intensity rather than duration alone. Together, these visual patterns substantiate the statistical improvements observed when adding category and hardware factors in the model-progression analysis (Table 2).
Figure 6. Distribution of average device power across laptops (C2–C5). Hardware-specific differences align with the coefficients in the duration-by-computer model ($R^{2}=0.960$), ranging from roughly 9 W for the most efficient to 12 W for the least efficient device.
[…]
Figure 8. Distribution of average device power across user flows. Power intensity varies by task type, confirming that flow-specific coefficients capture systematic rather than random variance.
5.2.2. Baseline and Hardware Effects
The global model ($E\!=\!10.15\times t$) achieved $R^{2}=0.946$, corresponding to an average effective power of 10.15 W. Introducing computer-specific coefficients increased the fit to $R^{2}=0.960$, revealing systematic power differences across devices (9.17–12.43 W). This reflects both CPU generation and OEM design variations.
5.2.3. Category and Interaction Effects.
Adding category-specific power factors improved the model to $R^{2}=0.973$, with typical fitted powers ranging from 8.4 W for navigation to 14.0 W for news tasks. Including detailed flow-level distinctions further increased $R^{2}$ to 0.978. When hardware and category were combined ($C(\text{category})\!:\!C(\text{laptop})$), the model explained $98.9\,\%$ of variance, and the full interaction model across flow and computer reached $R^{2}=0.994$. These results show that nearly all residual variance stems from identifiable and systematic differences in hardware and interaction type, rather than random error.
Table 1. Progression of OLS models explaining measured energy ($E$) from duration ($t$) and contextual factors (with intercept = 0).
Model Formula $R^{2}$
Global best-fit $E\sim t+0$ 0.946
+ Hardware (laptop) $E\sim t\!:\!C(\text{laptop})-1$ 0.960
+ Category $E\sim t\!:\!C(\text{category})-1$ 0.973
+ Flow $E\sim t\!:\!C(\text{flow})-1$ 0.978
Category $\times$ Hardware $E\sim t\!:\!C(\text{category})\!:\!C(\text{laptop})-1$ 0.989
Flow $\times$ Hardware $E\sim t\!:\!C(\text{flow})\!:\!C(\text{laptop})-1$ 0.994
5.2.4. Interpretation.
The progression demonstrates that hardware and web-category parameters substantially increase both precision and accuracy of energy estimation. The resulting context-aware model nearly reproduces measured values ($R^{2}=0.994$), confirming that the fixed-power assumptions of current reporting frameworks cannot capture the heterogeneity of realistic web-service energy use.
6. Discussion
The results show that deviations between modeled and measured energy are not random but scale systematically with interaction length, revealing a structural bias in the constant-power assumption of existing reporting models. The nearly perfect linear relation between Digst error and task duration confirms that this bias arises from the model formulation rather than from random measurement noise.
6.1. Improving Accuracy of Reporting Frameworks
The analysis in Section 5.2 showed that model deviations are systematic rather than random, driven primarily by reproducible differences in device power and service category. While incorporating hardware parameters can explain nearly all variance ($R^{2}=0.994$), this level of granularity is impractical for sustainability reporting. A more feasible improvement lies in introducing a limited set of representative service categories —such as shopping, booking, navigation, and news— which capture much of the same systematic variation ($R^{2}=0.973$). This categorical differentiation could substantially increase model precision while preserving the simplicity and comparability that reporting frameworks require. Such category grouping could therefore enhance the transparency and fairness of sustainability reporting without introducing additional measurement burden or model complexity.
6.2. Toward Reproducible and Transparent Sustainability Assessment
Digital sustainability models should incorporate reproducibility standards and sensitivity analyses that explicitly address hardware and user variance. Transparent reporting of assumptions, parameter sources, and version-controlled model dependencies would enable consistent comparisons across services and over time. The open repository accompanying this paper provides all datasets and scripts necessary to reproduce the results, including detailed software version information to ensure full experimental repeatability.
6.3. End-User Devices
Our measured laptop powers (9–13 W) align closely with modeled estimates in Malmodin’s long-term household measurements (malmodin2023, 19), which report average powers between 6 W and 12 W with short-term fluctuations from 3 W to 25 W. In addition, Malmodin’s results align well with the 10 W average derived in our best-fit model, which is realistic for browser-based workloads and representative of typical laptop energy behavior.
In 2013, Schien et al. (Schien2013, 27) observed average power dissipation of a laptop to be 32–37 W [^3] In the supplementary material, the energy consumption for 10 minutes of reading text and video watching was reported as 19.15 kJ and 22.01 kJ, respectively., and they found no statistically significant variation from idle power when browsing text. This does not align with our results. For comparison, see Table 2, which lists the Energy Star measured idle power $P_{idle}$ (hp_elitebook_830g6_2020, 17, 5, 6, 4), the minimum observed power during execution, and the best-fit power average. The best-fit power average is significantly higher than the Energy Star idle power and, in some cases, twice as high.
Table 2. For each laptop, the Energy Star $P_{idle}$, minimum observed power during execution, and the best-fit power dissipation average across all user flows.
No. Laptop $\mathbf{P}_{\mathbf{Idle}}$ Min Best-fit P
1 Dell Latitude 7490 6.53 W 9.74 W
2 HP EliteBook 830 G6 4.88 W 7.30 W 9.56 W
3 Dell Latitude E5250 6.32 W 6.74 W 9.17 W
4 Dell Latitude E5550 5.12 W 10.92 W 12.43 W
Malmodin also noted that Apple’s M-series architectures achieve averages below 6 W, and subsequent silicon generations, such as the M4 and newer Intel hybrid designs, are expected to further reduce typical laptop power dissipation. Thus, as hardware efficiency continues to improve, periodic recalibration of model parameters will be necessary to ensure that reported end-user power values remain representative of actual device behavior.
While our measured laptop powers (9–13 W) correspond well to these empirically grounded estimates, they are markedly lower than the fixed constants used in current sustainability reporting frameworks. For instance, the Danish Agency of Digital Government’s Digst model assumes 22 W, and the DIMPACT framework applies 15 W for browser-based scenarios. These simplified constants serve comparability purposes but appear conservative relative to modern laptop behavior, leading to systematic overestimation of end-user energy use in typical web interactions. Because both under- and overestimation of average power compound proportionally with interaction duration, any mismatch between assumed and actual device power introduces systematic rather than random error in total energy estimates. This observation suggests a need to explore alternative formulations that decouple energy estimation from pure duration, for instance by distinguishing a time-dependent baseline component from an activity-dependent component that captures variations in user interaction intensity.
6.4. Broader Context
Beyond laptop-based web interactions, similar mechanisms shape the energy profiles of other end-user workloads such as streaming and video communication. While this study focused on browser-based interactions, similar principles apply to other classes of end-user workloads, including streaming and video communication. Herglotz et al. (Herglotz2025, 16) demonstrated that energy measurement of video workloads requires precise synchronization between playback and power sampling. Our observations for the TV2 flow, which contained embedded autoplay videos, show markedly higher average power and total energy consumption than non-video pages, consistent with the controlled findings of Herglotz et al. (Herglotz2025, 16). Whereas their work isolated and synchronized video playback under laboratory conditions, our setup captured realistic end-user browsing sessions in which short videos appeared as part of ordinary news-site activity. Together, these perspectives confirm that even brief embedded videos can dominate device-level energy use, highlighting the need for category-specific modeling in sustainability reporting.
7. Limitations and Validity Considerations
To evaluate the robustness and scope of our findings, we examined potential threats to validity related to model structure, participant behavior, and experimental coverage.
Figure 9. Average measured energy across repeated runs (1–3). Energy decreased from Run 1 to Run 2 and remained stable thereafter, indicating a short learning effect without fatigue. Error bars represent standard deviations across participants.
Table 3. Number of distinct participants per flow–computer combination, showing balanced coverage across conditions.
Flow ID C2 C3 C4 C5
Aa 6 5 4 4
Ab 7 5 5 5
Ba 6 5 5 5
Bb 7 4 4 5
Ca 5 5 5 5
Cb 6 5 5 5
Da 6 5 5 5
Db 6 5 5 5
7.1. Model Structure and Bias
The proportional $P\times t$ formulation used in simplified reporting models inherently amplifies any misestimation of average device power. If the assumed power value is too high or too low, the resulting energy estimates accumulate a systematic bias proportional to task duration, as observed in Section 5.1. Consequently, even small parameter errors can produce large deviations for long-running interactions, limiting the external validity of constant-power approximations.
7.2. Participant Learning and Fatigue
Repeated participation may influence execution efficiency and thereby measured energy use. To mitigate such order effects, both the sequence of user flows and the allocation of laptops were randomized under constraints that ensured full coverage. Average energy use across all userflows declined from the first run (1453 J) to the second run (905J ) and then stabilized at (836 J) in the third run, see Figure 9 for the per userflow energy consumption. This pattern indicates a brief learning phase but no subsequent fatigue or performance drift. Because the learning effect was consistent across participants and evenly distributed through randomization, its influence on model validity is considered minimal.
7.3. Coverage and Inter-User Variance.
The experiment crossed eight user flows with four laptop platforms, resulting in a balanced dataset in which each flow–computer combination was represented by at least four and up to seven distinct participants (Table 3). This design ensured that device- and category-level effects were not driven by single-user behavior. Analysis of participant-level random effects in the mixed model (Section 5.2) yielded an intraclass correlation coefficient (ICC) of 0.012, indicating that only 1.2 % of total variance originated from between-participant differences. Thus, inter-user variability was negligible, and statistical power was sufficient for stable estimation of category- and device-level effects.
7.4. External Validity
The study focused on browser-based, non-streaming web services executed on four laptop platforms representative of common Intel generations and OEM designs. This scope ensured comparability across devices and categories but limits the generality of the findings in three respects. First, the tested service categories—shopping, booking, navigation, and news—represent only a subset of web interactions; extending the analysis to other categories such as streaming, productivity, or AI-assisted services would clarify the broader applicability of the results. Second, the dataset covers four laptop models, providing internal consistency but limited diversity across architectures. Expanding the hardware base to include a wider range of contemporary Energy Star–rated laptops would strengthen the trustworthiness of the derived power parameters. Third, the focus on laptops excludes other end-user devices such as mobile phones, tablets, and high-performance desktops, which may differ in baseline power and dynamic scaling behavior. All interactions were performed on live websites under real network conditions, ensuring ecological validity for laptop-based browsing. Consequently, the results are representative of realistic end-user conditions within this scope, while broader generalization to additional service types, hardware generations, and device classes should be examined in future studies.
8. Conclusion & Future Work
This study show that the constant-power assumptions used in digital sustainability reporting frameworks such as Digst and DIMPACT may systematically overestimate end-user device energy use. Empirical measurements of realistic browser-based interactions revealed typical laptop powers between 9 W and 13 W—substantially lower than the 22 W and 15 W values assumed in these models. As a result, model bias increases proportionally with task duration, producing structural rather than random error.
Assuming a near-perfect constant-power exists, we found that intorducing hardware and service-category parameters increases model precision markedly. This showed that most variance originates from reproducible, context-dependent differences in devices and workloads. These findings suggest that sustainability reporting can remain simple yet become substantially more accurate by adopting, e.g., category-specific power parameters.
All data, analysis scripts, and software versions are openly provided, enabling full reproducibility and future extensions of this work. Together, these results establish a methodological foundation for transparent, empirically grounded models of end-user device energy use, supporting more credible and comparable digital sustainability reporting.
Future work should aim to enhance the accuracy, transparency, and representativeness of end-user device energy models. First, expanding the hardware base to include a broader selection of contemporary laptops would improve the trustworthiness of derived parameters. Second, building on the evidence that constant-power assumptions systematically overestimate device energy, new models should explore ways to reduce systematic error by relating energy use to task context, interaction type, and hardware characteristics.
Although this study focused on browser-based interactions, it may be beneficial to explore whether the same principles could be extended to other software categories—such as streaming, productivity, and AI-assisted applications—to examine whether similar proportional relationships hold. Further research is required to refine category definitions and develop standardized, fair user flows to support reproducible end-user device energy assessments.
Acknowledgments
This work was supported by the Independent Research Fund Denmark Project no. 2102-00281B. We thank participants of the Symposium on Sustainable IT Systems for their contributions as participants and valuable inputs and especially Jens Malmodin for enlightening correspondence. We would like to thank Atea A/S for the donation of test computers. The authors acknowledge the use of generative AI for assistance on restructiring of text, language refinement, and consistency checks.